OpenStudy (anonymous):

how do you factor 4x^2-4 completely

OpenStudy (callisto):

4x^2-4 = 4(x^2-1) =4(x+1)(x-1)

OpenStudy (tommo_lcfc):

Take out a common factor of 4. 4(x^2 -1). Recognise that (x^2-1) is a difference of two squares. 4(x-1)(x+1).

OpenStudy (anonymous):

thank you both lol

OpenStudy (callisto):

welcome

OpenStudy (tommo_lcfc):

I explain things, @Callisto doesn't! That was a compliment to me, not an insult to her!

OpenStudy (tommo_lcfc):

Anyway, you can come to us any time. I think that me and @Callisto are a team now!

OpenStudy (callisto):

sorry, sometimes workings mean explanations to me. Just ask if you don't understand. people here are willing to guide you step by step :)

OpenStudy (anonymous):

lol well thank you and im gonna need it im sure lol math is my worst subject lol its much appreciated :)

OpenStudy (callisto):

math was also my worst subject :P

OpenStudy (anonymous):

lol well looks like you got it all figured ut now lol

OpenStudy (tommo_lcfc):

I love maths! I am at university doing straight maths :)

OpenStudy (anonymous):

lol right on, i wouldnt mind it if i actually knew what i was doin :)

OpenStudy (anonymous):

how would i factor 4x^2-4x+2xy-2...

OpenStudy (callisto):

take out the common factor 4x^2-4x+2xy-2 = 2(2x^2 -2x + xy-1)

OpenStudy (anonymous):

alright

OpenStudy (callisto):

because there is a y in the third term, you cannot take out common factor and group that again.. :(

OpenStudy (tommo_lcfc):

Could you take 2x out of the first two terms and get 2(2x(x-1)+xy-1) or is that not allowed?

OpenStudy (tommo_lcfc):

It can't be factorised any further after that.

OpenStudy (callisto):

there is no use to do so, i think?

OpenStudy (tommo_lcfc):

Probably not, if you say so.

OpenStudy (anonymous):

lol well either way the answer worked, so you were right lol :) thanks.

OpenStudy (callisto):

welcome :)